Current through branch a b

1
Expert's answer
2021-12-06T16:05:32-0500

Kindly note:


The information given in the question is not sufficient to solve the question.


Adjusting the conditions of the question to:


Find the current through the branch AB in the circuit diagram below.






Solution:





In loop XYBA:

"3I_2-1I_1=0\\I_1=3I_2"

In loop WZBA

"4(I_2+I_{ab})-2(I_1-I_{ab})=0"

"6I_{ab}=2I_1-4I_2 \\6I_{ab}=2(3I_2)-4I_2 \\I_2=3I_{ab}"


In loop ZQPY

"50-[3I_2+4(I_2+I_{ab})]=0\\50=7I_2+4I_{ab}"

"=7(3I_{ab})+4I_{ab} \\=25I_{ab}"


"I_{ab}=2A"